A sample of size 3 is taken from a population of the number of students in each class of Grade 9 students listed as 25, 30, 35, and 40.



Compute for the mean of the sampling distribution of the sample means.

1
Expert's answer
2022-03-16T05:52:45-0400

There are totally "{4 \choose 3}=4" samples of size 3. They are listed below

{25,30,35}; (25,30,40}; {25,35,40}; {30,35,40}

the mean of the sampling distribution of the sample means is "{\frac {{\frac {25+30+35} 3}+{\frac {25+30+40} 3}+{\frac {25+35+40} 3}+{\frac {30+35+40} 3}} 4}=32.5"